# Bound States at Semiconductor–Mott Insulator Interfaces

### v2

@@ III.1 Minimal Interface Perturbation
From an experimental standpoint, the selection of materials for the Mott insulator and the semiconductor will determine the interface perturbation $\Delta V$. Consequently, it is crucial to comprehend the requisite strength of the interface perturbation to establish a bound state. Interestingly, such a state is discovered to exist irrespective of whether $\Delta V$ is attractive or repulsive, provided it surpasses a specific threshold.
In order to calculate this threshold, we need to specify the allowed energy of a bound states. This is determined by the ansatz Eq. 16, which in the case $\Delta V>0$, is comprised of $r_{a}=r_{+}$ and $s_{b}=s_{-}$. To satisfy the condition $|s_{b}|=|s_{-}|<1$ any valid solution requires an energy $E>B=2T/Z-T_{\mathbf{k}}^{\|}+V$, ensuring that it lies energetically above the semiconducting band. To fulfill $|r_{+}|>1$, the upper band edges of the lower $A_{L}$ and of the upper $A_{H}$ Hubbard band are relevant:
@@
Secondly, for a semiconducting band edge situated between the center of the Mott gap and the upper Hubbard band $U/2<B<A_{H}$, the ansatz implies $E>A_{H}$. Consequently, Eq. 17 at $E=A_{H}$ yields:
@@
Thirdly, if the semiconducting band edge is between the lower Hubbard band and the middle of the Mott band gap $A_{L}<B<U/2$, any bound state has $B<E<U/2$. The threshold is found as $\Delta V_{\mathrm{min,B}}$.
@@
The required strength of the interface perturbation $\Delta V_{\mathrm{min}}$ depends on the band alignment. It is depicted in Fig. 3 (sold line) as a function of this band alignment, parametrized by the on-site potential $V$ for given hopping $T$. The approximate dependence of $T$, $U$, and $V$ is given. It decreases linearly with decreasing distance $V$ from the lower Hubbard band; is nearly constant in the lower half of the Mott gap; it goes up at $V=U/2$ and decreases from there linearly; before being constant again.
Figure 3: Minimally needed interface perturbation $\Delta V$ in the positive (solid) and negative (dashed) case as a function of the on-site potential $V$. The hopping strength $T=0.4U$ is fixed. The insets give the approximate formula for the minimal interface perturbation for $\Delta V>0$. The red dotted line gives the point symmetric rotated and shifted line of the positive interface perturbation onto the negative one.
@@
The minimum required interface perturbation can be calculated using the same approach as before, and is shown in Fig. 3 (dashed line). The solutions for attractive and repulsive $\Delta V$ seem to be related (at least approximately) via an inflection point inside the Hubbard gap. To illustrate this, the red dotted line in Fig. 3 is obtained by a shift plus an inflection from the solid line. It nearly matches the dashed one, with the differences coming from the different band edges that need to be considered in the two different cases.
@@ III.2 Decay Constant
The decay on both sides, $e^{-\kappa|\mu|}$, is characterized by an inverse length scale
$\kappa_{\mathrm{Mott}}$ $=\mathrm{arcosh}\left(\absolutevalue{\frac{Z}{2T}\left[\frac{E(U-E)}{E-U/2}-T_{\bf k}^{\|}\right]}\right),$ (23)
$\kappa_{\mathrm{semi}}$ $=\mathrm{arcosh}\left(\absolutevalue{\frac{Z}{2T}\left[V-E-T_{\bf k}^{\|}\right]}\right).$
@@
The decay constant for the semiconducting side is illustrated in Fig. 4. Notably for small $T$, there exists a range of interface perturbations without any solutions (e.g., for $T=0.2U$, from $\Delta V/U\approx 0.5-1$).
Figure 4: The decay constant in the unpolarized semiconducting half-space is presented for various on-site potentials $V$ and hopping strengths $T$ across the interface perturbation $\Delta V$. Black symbols correspond to $V=1.1U$, blue symbols to $V=1.5U$, and red symbols to $V=0.2U$. Different shapes represent distinct hopping strengths: circles for $T=0.2U$, triangles for $T=0.5U$, and squares for $T=0.8U$.
Figure 5: The decay constant in the unpolarized Mott half-space is shown for different on-site potentials $V$ and hopping strengths $T$ over the interface perturbation $\Delta V$. Black symbols correspond to $V=1.1U$, blue to $V=1.5U$ and red to $V=0.2U$. The different symbol shapes depict different hopping strengths, circles for $T=0.2$, triangles for $T=0.5U$ and squares for $T=0.8U$.
@@
Fig. 5 illustrates the same information for the Mott side. Again, for small hopping parameters (circles in Fig. 5), there is a range of $\Delta V$ values where no bound-state solution exists. In contrast to the semiconducting side, the decay constant in the Mott insulator may display non-monotonic behavior: When the on-site potential $V$ is close to the lower Hubbard band, a discontinuity occurs where $\kappa_{\mathrm{Mott}}$ sharply rises to a large value before abruptly decreasing close to zero. Increasing the interface perturbation for such a semiconducting band causes the energy of the solution to transition from being between the Hubbard bands to residing above the upper band, resulting in the observed discontinuity.
@@
The rate of decay, whether it occurs more rapidly in the Mott or the semiconducting side, depends on both the on-site potential and the hopping strength. Fig. 6 displays both decay constants for an on-site potential close to the lower Hubbard band and another above the upper one, both for a small hopping strength. In the case of the larger on-site potential, the decay constant in the Mott is greater than in the semiconductor, indicating faster decay in the Mott. However, for the smaller on-site potential near the lower Hubbard band, there is a reversal as the interface perturbation increases. For small interface perturbations, the decay is faster in the Mott, but as the interface perturbation reaches a certain magnitude, the decay becomes slower than in the semiconductor.
@@
Figure 6: Decay constant for both sides with $V=0.2U$ (black) and $V=1.5U$ (blue dashed) with $T=0.2U$ in the unpolarized case. The semi lines are on top of each other. Note the gap in the black curves.
@@
Figure 7: Decay constant for a negatice interface perturbation $\Delta V$ in the unpolarized case. Circles give the $\kappa_{\mathrm{Mott}}$ and triangles give $\kappa_{\mathrm{semi}}$. The different colors depict different on-site potentials $V$. Hopping strength $T=0.2U$.
@@
Experimental measurements at the $\mathrm{LaTiO}_{3}/\mathrm{SrTiO}_{3}$ system show a tunnelling strength of $T=0.3\mathrm{eV}$ with a Coulomb repulsion of $U=6-20T$, such that realistic values are in the $T=0.16-0.05U$ regime. This matches the here used value of $T=0.2U$. The bound states found here have dispersion in the parallel direction. It is likely that all these subbands are partially filled leading to a metallic state parallel to the interface as it was measured in [1, 15, 32].
@@ IV Unpolarized Mott region between two semiconductors
Utilizing the same methodology, we can also compute bound states in a system with two interfaces, where an unpolarized Mott insulator is positioned between two semi-infinite semiconductors, similar to a quantum well. Notably, there is no pre-requisite for the semiconductors to be identical; they can have distinct on-site potentials $V$ and interface perturbations $\Delta V$. The Mott region now spans $-a\leq\mu\leq a$, leading to the following ansatz:
$p_{n}^{0}=\begin{cases}As_{1}^{\mu}&\mu<-a\\ Br_{+}^{\mu}+Cr_{-}^{\mu}&-a\leq\mu\leq a\\ Bs_{2}^{\mu}&a<\mu\end{cases}$ (24)
Here, $|s_{1}|>1$ and $|s_{2}|<1$, with no requirement on the absolute value in the Mott region. The four boundary conditions now yield two conditions relating the amplitudes:
11 11 $=\frac{2(E-U)}{2E-U}\frac{s_{1}^{a+1}}{A}\left(Br_{+}^{-a-1}+Cr_{+}^{a+1}\right),$ (25)
11 11 $=\frac{2(E-U)}{2E-U}\frac{s_{2}^{-a-1}}{D}\left(Br_{+}^{a+1}+Cr_{+}^{-a-1}\right).$
@@
In the symmetric case, where the left and right semiconductors are the same, $s_{2}=s_{\bar{1}}$ holds to satisfy the condition of absolute values (with the bar denoting the opposite index).
@@
Now, there exist even ($+$) solutions (with $B=C$, $A=D$) and odd (--) solutions (with $B=-C$, $A=-D$). With this, the boundary conditions for the coefficients and the defining equation for bound states in this system are given by:
11 11 $=\frac{2(E-U)}{2E-U}\frac{B}{A}\left(r_{+}^{-a-1}\pm r_{+}^{a+1}\right)s_{1}^{a+1},$ (27)
$\frac{\Delta VZ}{T}$ $=\frac{r_{+}^{a}\pm r_{+}^{-a}}{r_{+}^{a+1}\pm r_{+}^{-a-1}}-s_{1}.$
Here, $+$ denotes the even parity case, while -- indicates the odd one. Choosing $s_{1}=s_{\pm}$ alters the allowed energy range of the solutions to meet the requirement $|s_{1}|>1$; $s_{+}$ yields allowed energies above the semiconducting band, while $s_{-}$ yields those below.
IV.1 Standing-wave solutions
@@
Figure 8: Defining equation with $\Delta V=-0.6U$, $T=0.4U$ and $V=1.5U$ in the zero parallel momentum sector and $a=6$. The crossings of the solid and dashed line with the dotted one give the energies for the even and odd parity solutions, respectively.
@@
Disregarding any interface perturbation, the energy of the solutions must fall into one of the two Mott bands (see Fig. 8). Defining the Mott region as $\mu\in[-a,a]$, i.e. with $12a+$ 1 Mott lattice sites, we obtain $22a+$ 2 even-parity solutions from the defining equation. These are evenly distributed into $a+1$ solutions in both the lower and upper Hubbard bands. For odd parity, there are $2a$ solutions, equally divided into $a$ solutions in the lower and upper Hubbard bands. All of them have quasi-particle density leaking from the Mott into the semiconductors.
@@
The fulfillment of the absolute value condition depends solely on the choice of $s_{1}$ and the potential $V$. A change in the potential has a minimal impact on the solutions, slightly lowering the energy (analogous to a potential well), as well as on the leaking. For solutions in the lower Hubbard band, the leakage is nearly independent of the on-site potential.
@@
The solutions manifest as standing waves with $|r_{\pm}|=1$ in the Mott region. In both the lower and upper Hubbard bands, the solutions have energies $E_{0},E_{1},...,E_{2a}$, where even energies $E_{0}$, $E_{2}$, etc., correspond to even parity solutions, and odd energies $E_{1}$, $E_{3}$, and so forth, correspond to odd parity ones. The number of zeros of $p_{\mu}^{0}$ increases with energy from 0 to $2a$.
@@
Enumerating only the even energy solutions as $E^{\mathrm{even}}_{0},E^{\mathrm{even}}_{1},\ldots E^{\mathrm{even}}_{a}$, the corresponding quasi-particle probability distributions $|p_{\mu}^{0}(E_{n}^{\mathrm{even}})|^{2}$ have $2(\frac{a}{2}-|\frac{a}{2}-n|)$ nodes. For $a=6$, this results in the sequence $0-2-4-6-4-2-0$. Performing the same analysis for the odd energy solutions $E^{\mathrm{even}}_{0},E^{\mathrm{even}}_{1},\ldots E^{\mathrm{even}}_{a-1}$, this yields $12(\frac{a}{2}-|\frac{a}{2}-n|)+$ 1 nodes ($1-3-5-5-3-1$ for $a=6$). In the lower Hubbard band the quasi-particle probability distribution is symmetric between states with the same number of nodes. In the upper Hubbard band, energies denoted by $E_{Un}$ ,the symmetry of the probability distribution is broken. As the energy of the solution increases, more of the probability becomes localized at the interfaces; with this the amount of leaked density increases as well.
Lowering the on-site potential $V$ such that the left band edge of the semiconducting band is below the lower Hubbard band results in the absence of $s_{1}=s_{-}$ solutions. Instead, they are replaced by $s_{1}=s_{+}$ solutions. For intermediate potentials with bands completely within the Mott gap, the lower Hubbard solutions are given by $s_{1}=s_{-}$, while the upper band ones are given by $s_{1}=s_{+}$. The energy of the solutions, as well as their $p_{\mu}^{0}$, remains unchanged when changing $s_{1}$ from $s_{-}$ to $s_{+}$; thus, bound state solutions persist within this heterostructure for all $V$.
@@
As in the single interface case, the bound states display a dispersion in the parallel direction leading to subbands. These subbands are partially filled and as there is quasi particle weight at the interface the bound states exhibit metallic behavior there [1, 32, 15].

### v3

@@ III.1 Minimal Interface Perturbation
From an experimental standpoint, the selection of materials for the Mott insulator and the semiconductor will determine the interface perturbation $\Delta V$, the hopping strength $T$ as well as the band offset $V$. Consequently, it is crucial to comprehend the requisite strength of the interface perturbation to establish a bound state. Without this, the interface only supports interface resonances or, in case the band overlap energetically, quasi-particle wave functions covering both half-spaces. Interestingly, such a state is discovered to exist irrespective of whether $\Delta V$ is attractive or repulsive, provided it surpasses a specific threshold. While experimentally the choice of materials fixes the parameters, in our theoretical study it is instructive to tune the band offset $V$. As bound states have an energy outside of the Mott and semiconducting band, the respective band edges serve as a natural boundary for the allowed energies. For this, we need to distinguish between an attractive and repulsive perturbation.
In the $\Delta V>0$ case, bound states are comprised of $r_{a}=r_{+}$ and $s_{b}=s_{-}$. To satisfy the condition $|s_{b}|=|s_{-}|<1$ any valid solution requires an energy $E>B=2T/Z-T_{\mathbf{k}}^{\|}+V$, ensuring that it lies energetically above the semiconducting band. To fulfill $|r_{+}|>1$, the upper band edges of the lower $A_{L}$ and of the upper $A_{H}$ Hubbard band are relevant:
@@
Secondly, for a semiconducting band edge situated between the center of the Mott gap and the upper Hubbard band $U/2<B<A_{H}$, the ansatz implies $E>A_{H}$. This is the case in Ref. [55]. Consequently, Eq. 19 at $E=A_{H}$ yields:
@@
Thirdly, if the semiconducting band edge is between the lower Hubbard band and the middle of the Mott band gap $A_{L}<B<U/2$, any bound state has $B<E<U/2$. The threshold is found as $\Delta V_{\mathrm{min,B}}$. This is the case in Ref. [71].
@@
The required strength of the interface perturbation $\Delta V_{\mathrm{min}}$ depends on the band alignment. It is depicted in Fig. 3 (sold line) as a function of this band alignment, parametrized by the potential offset $V$ for given hopping $T$. The approximate dependence of $T$, $U$, and $V$ is given. It decreases linearly with decreasing distance $V$ from the lower Hubbard band; is nearly constant in the lower half of the Mott gap; it goes up at $V=U/2$ and decreases from there linearly; before being constant again. For an existing interface, the parameter combination is given by the materials such that this system would be a single point on either the solid or dashed line (two examples are marked). By a gate voltage applied to only one half-space, the band offset $V$ may be shifted, and the minimally needed strength moves accordingly.
Figure 3: Minimally needed interface perturbation $\Delta V_{\mathrm{min}}$ in the positive (solid) and negative (dashed) case as a function of the on-site potential offset $V$. The hopping strength $T=0.4U$ is fixed. The annotations give the approximate formula in the $\Delta V>0$ case. The vertical lines give the band offsets for $\mathrm{SrTiO}_{3}$ interface with $\mathrm{SmTiO}_{3}$ and $\mathrm{LaVO}_{3}$, respectively.
@@
The minimum required interface perturbation can be calculated using the same approach as before, and is shown in Fig. 3 (dashed line). The solutions for attractive and repulsive $\Delta V$ seem to be related (at least approximately) via an inflection point inside the Hubbard gap.
@@ III.2 Decay Constant
The decay on both sides, $e^{-\kappa|\mu|}$, is characterized by an inverse length scale:
[…]
@@
The decay constant for the semiconducting side is illustrated in Fig. 4. Notably for small but realistic $T$, there exists a range of interface perturbations without any solutions (e.g., for $T=0.2U$, from $\Delta V/U\approx 0.5-1$) for a band offset close to the lower Hubbard band $V=0.2U$, which is close to the measured ones for $\mathrm{SrTiO}_{3}/\mathrm{SmTiO}_{3}$ and $\mathrm{SrTiO}_{3}/\mathrm{GdTiO}_{3}$ [71].
Figure 4: The decay constant $\kappa_{\text{semi}}$ in the unpolarized semiconducting half-space is shown as a function of the interface perturbation for different band offsets $V$. The hopping strength is fixed at $T=0.2U$. Note that gap in the $0.2U$ data.
Figure 5: The decay constant $\kappa_{\text{Mott}}$ in the unpolarized Mott half-space is shown for different band offsets $V$ as a function of the interface perturbation $\Delta V$. The hopping strength is fixed at $T=0.2U$.
@@
Fig. 5 illustrates the same information for the Mott side. Again, for realistic hopping parameters there is a range of $\Delta V$ values where no bound-state solution exists. In contrast to the semiconducting side, the decay constant in the Mott insulator may display non-monotonic behavior: When the band offset $V$ is close to the lower Hubbard band, a discontinuity occurs where $\kappa_{\mathrm{Mott}}$ sharply rises to a large value before abruptly decreasing close to zero. Increasing the interface perturbation for such a semiconducting band causes the energy of the solution to transition from being between the Hubbard bands to residing above the upper band, resulting in the observed discontinuity.
@@
The rate of decay, whether it occurs more rapidly in the Mott or the semiconducting side, depends on both the band offset and the hopping strength. Fig. 6 displays both decay constants for a band offset close to the lower Hubbard band and another above the upper one, both for a small but realistic hopping strength. In the case of the larger on-site potential, the decay constant in the Mott is greater than in the semiconductor, indicating faster decay in the Mott. However, for the smaller on-site potential near the lower Hubbard band, there is a reversal as the interface perturbation increases. For small interface perturbations, the decay is faster in the Mott, but as the interface perturbation reaches a certain magnitude, the decay becomes slower than in the semiconductor.
@@
Figure 6: Decay constants $\kappa_{\mathrm{Mott}}$$\kappa_{\mathrm{semi}}$ for both sides with $V=0.2U$ (red) and $V=1.5U$ (blue dashed) with $T=0.2U$ in the unpolarized case. Note the gap in the red square and diamond curves.
@@
Figure 7: Decay constants $\kappa_{\mathrm{Mott}}$$\kappa_{\mathrm{semi}}$ for a negative interface perturbation $\Delta V$ in the unpolarized case. The different colors depict different on-site potentials $V$. Hopping strength $T=0.2U$.
@@
Experimental measurements at the $\mathrm{LaTiO}_{3}/\mathrm{SrTiO}_{3}$ system show a tunnelling strength of $T=0.3\mathrm{eV}$ with a Coulomb repulsion of U=620TU=6-20T, such that realistic values are in the T=0.160.05UT=0.16-0.05U regime. This matches the here used value of $T=0.2U$. The decay constant for these parameters yields bound states with a non-zero quasi-particle density extending a few layers around the interface, consistent with other works [10, 79]. The bound states found here have dispersion in the parallel direction. It is likely that all these subbands are partially filled leading to a metallic state parallel to the interface as it was measured in [1, 15, 32].
@@ IV Unpolarized Mott region between two semiconductors
Utilizing the same methodology, we can also compute bound states for the quasi-particle wave functions in a system with two interfaces, where an unpolarized Mott insulator is positioned between two semi-infinite semiconductors, similar to a quantum well. Notably, there is no pre-requisite for the semiconductors to be identical; they can have distinct band offsets $V$ and interface perturbations $\Delta V$. The Mott region now spans the site indices $a\leq\mu\leq a$, leading to the following ansatz:
$p_{n}^{0}=\begin{cases}As_{1}^{\mu}&\mu<-a,\\ Br_{+}^{\mu}+Cr_{-}^{\mu}&-a\leq\mu\leq a,\\ Ds_{2}^{\mu}&a<\mu.\end{cases}$ (26)
Here, $|s_{1}|>1$ and $|s_{2}|<1$, with no requirement on the absolute value in the Mott region. Remember that the $s_{i}$ give the electron wave function eigenmodes in the semiconductor while the $r_{i}$ give the quasi-particle eigenmodes in the Mott insulator. The four boundary conditions now yield two conditions relating the amplitudes:
$1$ $=\frac{2(E-U)}{2E-U}\frac{s_{1}^{a+1}}{A}\left(Br_{+}^{-a-1}+Cr_{+}^{a+1}\right),$ (27)
$1$ $=\frac{2(E-U)}{2E-U}\frac{s_{2}^{-a-1}}{D}\left(Br_{+}^{a+1}+Cr_{+}^{-a-1}\right).$
@@
In the symmetric case, where the left and right semiconductors are the same, $s_{2}=s_{\bar{1}}$ holds to satisfy the condition of absolute values (with the bar denoting the opposite index). This ensures that the bound state does decay inside both semiconducting half-spaces.
@@
Now, there exist even (++) solutions (with $B=C$, $A=D$) and odd (-) solutions (with B=CB=-C, A=DA=-D). With this, the boundary conditions for the coefficients and the defining equation for bound states in this system are given by:
$1$ $=\frac{2(E-U)}{2E-U}\frac{B}{A}\left(r_{+}^{-a-1}\pm r_{+}^{a+1}\right)s_{1}^{a+1},$ (29)
$\frac{\Delta VZ}{T}+s_{1}$ $=\frac{r_{+}^{a}\pm r_{+}^{-a}}{r_{+}^{a+1}\pm r_{+}^{-a-1}}.$ (30)
Here, ++ denotes the even parity case, while - indicates the odd one. The energy of the bound state physically fixes the mathematical ansatz for the $s_{i}$. $s_{1}=s_{+}$ yields energies greater than the semiconducting band, while $s_{-}$ yields those below. The second equation is the defining one for quasi-particle bound states in this double interface system, the first one normalizes the wave function amplitudes.
IV.1 Standing Wave Solutions
@@
Figure 8: Defining equation (27) with $\Delta V=-0.6U$, $T=0.4U$ and $V=1.5U$ in the zero parallel momentum sector and $a=6$. The crossings of the solid and dashed line with the dotted one give the energies for the even and odd parity solutions, respectively.
@@
Disregarding any interface perturbation, the bound state energy must fall into one of the two Mott bands (see Fig. 8). Defining the Mott region as $\mu\in[-a,a]$, i.e. with $2a+1$ Mott lattice sites, there are even and odd parity solutions for the quasi-particle wave function. This is well known from the standard quantum well. Mathematically, we obtain $2a+2$ even-parity solutions for the wave function, which are evenly distributed into $a+1$ solutions in both the lower and upper Hubbard bands. For odd parity, there are $2a$ solutions, equally divided into $a$ solutions in the lower and upper Hubbard bands. All of them have quasi-particle density leaking from the Mott into the semiconductors.
@@
The physical condition of the wave function decay away from the interfaces as well as the band offset $V$ set the mathematical ansatz for $s_{1}$. A change in the offset has a minimal impact on the solutions, slightly lowering the energy (analogous to a potential well), as well as on the leaking. For solutions in the lower Hubbard band, the leakage is nearly independent of the band offset.
@@
The doublon and holon wave functions manifest as standing waves with $|e^{\pm i\kappa_{\mathrm{Mott}}\mu}|=1$ in the Mott region. As known from the quantum well, even and odd parity solutions alternate in energy. Hence, in both the lower and upper Hubbard bands, the standing waves have energies $E_{0}^{L},E_{1}^{L},\ldots,E_{2a}^{L}$, and $E_{0}^{U},E_{1}^{U},\ldots E_{2a}^{U}$ respectively, where even (odd) index energies correspond to even (odd) parity solutions. Additionally, and again analogous to the quantum well, the number of nodes of the holon wave function $p_{\mu}^{0}$ increases with energy from 0 to $2a$.
@@
More precisely, the corresponding quasi-particle probability distributions $|p_{\mu}^{0}(E_{n})|^{2}$ have $2(\frac{a}{2}-|\frac{a}{2}-\frac{n}{2}|)$ nodes for $n$ even. For $a=6$, this results in the sequence 02464200-2-4-6-4-2-0. Performing the same analysis for the odd index energy solutions yields $2(\frac{a}{2}-|\frac{a}{2}-\frac{n}{2}|)+1$ nodes, (1355311-3-5-5-3-1 for $a=6$). In the lower Hubbard band the quasi-particle probability distribution is the same between states with the same number of nodes independent of their energy. In the upper Hubbard band, this is not true anymore. As the energy of the solution increases, more of the probability becomes localized at the interfaces; thereby the amount of leaked density increases as well.
@@
Bound state solutions persist within this heterostructure for all $V$, independent of the band offset. As in the single-interface case, the bound states display a dispersion in the parallel direction, leading to subbands. These subbands are partially filled and, as there is quasi particle weight at the interface, the bound states exhibit metallic behavior there [1, 32, 15].